OpenStudy (anonymous):

red kangaroo jumps along at 40km/h At this rate how long would it take to jump 2km

OpenStudy (anonymous):

40 km = 1hr 2 km = (1*2)/40hrs = 1/20 hrs = (1/20)*60 min = 3 min = 3 * 60 sec = 180 sec

OpenStudy (anonymous):

don't forget to acknowledge those who help u with a medal or fan

OpenStudy (anonymous):

40 km =1 h 2 km = x x=(2 km* 1h)/40 km =2/40=1/20 h now 1 h = 60 min 1/20 h =x x=( 1/20 h * 60 min)/1 h= 60/20=3 min